OpenStudy (anonymous):

2-4|x+2|>= -6 Help please show steps!

OpenStudy (anonymous):

-4|x+2|>=-3 |x+2| <= 3/4 |x| <= -5/4 Should be right. It checks out when plugged in the equation. Someone verify that

OpenStudy (anonymous):

Take the 2 across then divide by -4 (flips ineq). Then use def of abs value (|x|<=1 means -1<=x<=1) and then minus 2 from all three parts of the double ineq. To get -4<=x<=0

OpenStudy (anonymous):

Damn iunno where that -3 came from. Ignore my answer lol